OpenStudy (anonymous):

Which of the following is equivalent to 5 2/ 5 divided by 10?

OpenStudy (anonymous):

@angel12310

OpenStudy (anonymous):

Which of the following is equivalent to 5 2/ 5 divided by 10?

OpenStudy (anonymous):

OH HI

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

Which of the following is equivalent to 5 2/ 5 divided by 10? 1 23/27 1/2 27/50 5 1/25

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

27/5 ÷ 10/1 = (27 × 1) / (10 × 5) = 27/50 = 27/50

OpenStudy (anonymous):

anymore

OpenStudy (anonymous):

THANKS, YES I HAVE MORE

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

by the way i love your pic

OpenStudy (anonymous):

colorful

OpenStudy (anonymous):

THANKS IS THAT YOU IN THE PICTURE

OpenStudy (anonymous):

no a singer

OpenStudy (anonymous):

well actress

OpenStudy (anonymous):

ok questions

OpenStudy (anonymous):

MATCH THE FOLLOWING: -1/6 - 2/3 -2/3 - (-5/6 ) -1/2 - (-1/3) -5/6 + 1/2 CHOICES: 1/6, -1/3, -5/6, -1/6

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

different one

OpenStudy (anonymous):

LIKE THE ANSWER

OpenStudy (anonymous):

MATCH THE FOLLOWING: -1/6 - 2/3 = -2/3 - (-5/6 ) = -1/2 - (-1/3) = -5/6 + 1/2 = CHOICES: 1/6, -1/3, -5/6, -1/6

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

YEH

OpenStudy (anonymous):

-1/6 - 2/3 =-5/6

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

-2/3 - (-5/6 )=1/6

OpenStudy (anonymous):

-1/2 - (-1/3)=-1/6

OpenStudy (anonymous):

-5/6 + 1/2 =-1/3

OpenStudy (anonymous):

there

OpenStudy (anonymous):

OK THANKS

OpenStudy (anonymous):

WAIT I HAVE MORE

OpenStudy (anonymous):

Renee took a long flight to visit her family that lives out of the country. She slept for the first quarter of the trip and watched movies for half of the remaining flight. If the total trip took 9 3/4 hours, how long did Renee watch movies? 3 21/32 3 1/24 3 5/12 5 2/3

OpenStudy (anonymous):

HELLO?ARE U THERE

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hello I have a ques.

OpenStudy (anonymous):

Add -4 3/8 + -5 3/4= -10 1/8 -8 1/8 -1 3/8 -3/8

OpenStudy (anonymous):

-35/8 + -23/4 = ((-35 × 4) + (-23 × 8)) / (8 × 4) = (-140 + -184) / 32 = -324/32 = -81/8 = -10 1/8

OpenStudy (anonymous):

@happyface7